OpenStudy (anonymous):

Solve the system. 5(x-1)=29-3(y+2) 4(x+2)-22=3(2-y) (x,y) = ( , ) Suppose to use the elimination or substitution method but i cant figure it out please help thx

OpenStudy (cwrw238):

you'll find its easier if you rearranged the equations into the form ax + by = c

OpenStudy (anonymous):

5x-5=29-3y+6 5x+3y=40 Right?

OpenStudy (cwrw238):

nearly its - 6 n first line

OpenStudy (anonymous):

4x+8-22=6-3y 4x+3y=20

OpenStudy (cwrw238):

it comes to 5x + 3y = 28

OpenStudy (anonymous):

ok so 5x+3y=28 4x+3y=20

OpenStudy (cwrw238):

right elimination method is quite easy just subtract and you;ve eliminated y

OpenStudy (anonymous):

5x+3y=28 -4x-3y=-20 -4x=8 x=-2?

OpenStudy (anonymous):

y=-4?

OpenStudy (cwrw238):

- and got x straight away

OpenStudy (anonymous):

What?

OpenStudy (cwrw238):

whys multiply by -1? you've got 3y in each equation 3y-3y = 0

OpenStudy (anonymous):

so x=8?

OpenStudy (cwrw238):

yes simple

OpenStudy (cwrw238):

now plug x=8 into any of the equations to get y

OpenStudy (anonymous):

y=4

OpenStudy (cwrw238):

no try again

OpenStudy (cwrw238):

wrong sign

OpenStudy (anonymous):

-4

OpenStudy (cwrw238):

4(8) + 3y = 20 3y = 20-32 = -12 y = -4

OpenStudy (anonymous):

thx :DDDDD